# Measure-Theoretic Time-Delay Embedding

### v1

@@ 4.2 First Examples: Noisy Chaotic Attractors
[…]
4.3 NOAA Sea Surface Temperature Reconstruction
We now consider the problem of reconstructing the NOAA Sea Surface Temperature (SST) from partial measurement data [41]. The SST dataset consists of weekly temperature measurements sampled at a geospatial resolution of $1\degree$; see the supplementary materials for a visualization. Our partial observation of the full SST dataset $\{\mathbf{z}(t_{i})\}\subseteq\mathbb{R}^{44219}$ consists of the temperature time-series $\{x(t_{i})\}\subseteq\mathbb{R}$ recorded at the location $(156\degree,40\degree)$. Our goal is to use the delay state corresponding to $\{x(t_{i})\}$ to learn a reconstruction map for the full state $\{\mathbf{z}(t_{i})\}$. A similar problem is considered in [12, 35].
@@ 4.3 NOAA Sea Surface Temperature Reconstruction
The dataset is partitioned into training and testing sets, where the training set contains 355 snapshots ($\approx 5$ years) and the testing set contains 1415 snapshots (27($\approx$ 27( ≈ 27 years). Based on the time-series $\{x(t_{i})\}$, we select a time delay of $\tau=12$ (weeks) and an embedding dimension $m=7$. To reduce the computational cost in learning, similar to [35, 12], we subtract off the temporal mean of the SST dataset and parameterize the full state of the system by the first $N_{\texttt{POD}}$ time-varying POD coefficients, $\{\alpha_{k}(t)\}_{k=1}^{N_{\texttt{POD}}}$, which are obtained via the method of snapshots; see [35, Section 3.1]. That is, we perform the model reduction
@@
For this problem, all training samples are normalized via an affine transformation, such that the $L^{\infty}$ norm of the data vectors is at most 1. We train the neural network parameterization $\mathcal{R}_{\theta}$ using both the pointwise ((13)) and measure-based ((14)) approaches. The pointwise approach seeks to directly enforce the relationship (17) through the mean-squared error, whereas the measure-based approach partitions the delay state into clusters and aims to push forward the empirical measure in each cluster into the corresponding measure in the POD space. We use a constrained k-means routine to evenly partition the delay state into $5$ clusters. A visualization of the 5 clusters can be found in the supplementary materials. We evaluate the performance of the learned reconstruction map by forecasting the time-varying POD coefficients $\{\alpha_{k}(t_{i})\}$ for the testing set, which is then used to reconstruct the full state $\{\mathbf{z}(t_{i})\}$ according to (16).
We note that the challenge of learning the reconstruction map $\mathcal{R}_{\theta}:\mathbb{R}^{7}\to\mathbb{R}^{200}$ is exacerbated by the sparsity of the available training data. Specifically, we aim to learn a $200$-dimensional map using only $355$ training examples. Due to this data sparsity, we utilize MMD based on the Gaussian kernel $k_{\text{g}}(x,y)=\exp(-\|x-y\|_{2}^{2}/2\sigma^{2})$ with $\sigma=3$. The choice of a relatively large $\sigma$ acts as a form of regularization, helping to mitigate overfitting when dealing with sparse samples [18].
@@
Table 3: Reconstruction mean squared error (MSE) for the SST dataset after various numbers of training iterations. The measure-based approach ((14)) is less prone to overfitting than the pointwise approach ((13)).
In Fig. 3, we visualize the pointwise and measure-based reconstructions of the SST example at testing weeks 425, 550, 675, and 800 after training both models for $25{,}000$ steps. One can observe that the measure-based results align more closely with the ground-truth snapshots, while the pointwise reconstructions exhibit numerous nonphysical oscillations.
Figure 3: Visual comparison of the pointwise ((13)) and measure-based ((14)) approaches to reconstructing the SST dataset at the testing weeks 425, 550, 675, and 800. The left column features the ground truth snapshot, the middle column shows the measure-based reconstruction, and the right column shows the pointwise reconstruction.
@@
Table 4: Testing reconstruction mean squared error (MSE) for the ERA5 wind speed dataset after various iterations during training. For each loss function, the experiment was repeated $5$ times to approximate the mean and standard deviation for the reconstruction error.
@@ 4.4 ERA5 Wind Field Reconstruction
Our final test reconstructs a portion of the ERA5 wind speed dataset using partial measurement data [24]. We consider the ERA5 wind dataset sampled at a geopotential height of Z500, restricted to latitudes between $-180\degree$ and $-105\degree$ and longitudes between $-90\degree$ and $-15\degree$. The dataset is coarsened in both space and time, with measurements taken on approximately a $4\degree$ spatial grid with dimensions $20\times 20$, and separated by $6$ hours in time. For training, we use $2\times 10^{3}$ randomly sampled snapshots of both the $u$ and $v$ wind speed components (longitudinal and latitudinal, respectively) between 2000 and 2007. For testing, we consider the subsequent $5\times 10^{3}$ consecutive snapshots.
@@
In (18), the time-delay and embedding dimension are heuristically chosen as $\tau=6$ (hours) and $m=4$. In (19), the indices $\{n_{k}\}_{k=1}^{30}$ correspond to 30 randomly sampled spatial locations at which we observe the longitudinal and latitudinal components of the wind speed. The full longitudinal wind speed vector is $\mathbf{u}(t_{i})=\left(u_{1}(t_{i}),\dots,u_{400}(t_{i})\right)\in\mathbb{R}^{400}$ representing values at the $20\times 20$ spatial grid, while the full latitudinal wind speed vector is $\mathbf{v}(t_{i})=\left(v_{1}(t_{i}),\dots,v_{400}(t_{i})\right)\in\mathbb{R}^{400}.$ The combined full state can then be written $\mathbf{z}(t_{i})=(\mathbf{u}(t_{i}),\mathbf{v}(t_{i}))\in\mathbb{R}^{800}$, which is precisely the quantity we seek to predict from the delay state $\mathbf{y}(t_{i})$; see (18).
We train models to approximate the state reconstruction map $\mathbf{y}(t_{i})\mapsto\mathbf{z}(t_{i})$ using both the pointwise reconstruction loss ((13)) and the measure-based reconstruction loss ((14)). Throughout, we normalize the training data by subtracting the temporal mean and ensuring that each feature has unit variance. Recall that we randomly sample $2\times 10^{3}$ times $t_{i}$ at which we observe both the delay state $\mathbf{y}(t_{i})$ and the full state $\mathbf{z}(t_{i})$, allowing us to form the paired data $\{(\mathbf{z}(t_{i}),\mathbf{y}(t_{i}))\}_{i=1}^{2000}.$ Thus, when learning the reconstruction map according to the pointwise loss ((13)), we directly enforce the relationship $\mathbf{y}(t_{i})\mapsto\mathbf{z}(t_{i})$, for $1\leq i\leq 2\times 10^{3}$, via the mean squared error.
In contrast, when learning the reconstruction map using the measure-based loss (14), we transform the paired data $\{(\mathbf{z}(t_{i}),\mathbf{y}(t_{i}))\}_{i=1}^{2000}$ into probability measures and aim to find a suitable pushforward map between the probability measures in the time-delayed state space and those in the full state space. We first apply a constrained k-means clustering algorithm to partition the time-delayed samples $\{\mathbf{y}(t_{i})\}_{i=1}^{2000}\subseteq\mathbb{R}^{240}$ into 20 distinct clusters, each containing 100 samples in $\mathbb{R}^{240}$. In doing so, we obtain 20 pairs of probability distributions. During training, we then learn a model that pushes forward each measure in the time-delay coordinate system to the corresponding measure in the full state space. We evaluate the measure-based approach (14) using both the Energy MMD loss (with a polynomial kernel $k_{\text{p}}$) and the Gaussian MMD loss (with a Gaussian kernel $k_{\text{g}}$ and $\sigma=25$).
@@
It is worth noting that choosing a relatively large bandwidth for the Gaussian MMD acts as additional regularization, helping to prevent overfitting to the locations of the training samples. For longer training times, the Gaussian MMD outperforms the Energy MMD in measure-based reconstruction. The reconstructions based on the Gaussian MMD are smoother than the pointwise reconstructions and more closely match the ground truth; a visualization can be found in the supplementary materials. In Fig. 4, we show the average spatial distribution of the reconstruction error for the Gaussian MMD and pointwise approaches.
@@ 5 Conclusion
In this work, we have introduced a significant advancement to the classical Takens’ embedding theorem by developing a measure-theoretic generalization. While the original theory focuses on the embedding property of individual time-delayed trajectories, our novel approach shifts the focus to probability distributions over the state space. Through Corollary 1, our main theoretical contribution, we have demonstrated that the embedding indeed occurs within the space of probability distributions under the pushforward action of the time-delay map. This breakthrough was made possible by integrating the classical Takens’ theorem with cutting-edge tools from optimal transport theory, which were pivotal in our analysis.
@@
Building on these theoretical foundations, we devised a measure-theoretic computational routine for learning the inverse embedding map, also known as the full-state reconstruction map. This innovative method enables the forecasting of an entire high-dimensional dynamical system state from the time lags of a single observable. Our approach involved partitioning the observed trajectory in time-delay coordinates using a k-means clustering routine, where each cluster represents discrete probability measures. During training, we ensured that the corresponding discrete measures in the reconstruction space matched the pushforward distributions of the measures in the time-delay coordinates. This training scheme represents a relaxation of classical pointwise matching, allowing for greater tolerance of pointwise errors in the final reconstruction based on the scale of each discrete measure in the reconstruction space.
While classical pointwise matching may yield more accurate results in the ideal noise-free and densely sampled data scenario, such conditions are rarely met in real-world applications. Our method shines in practical situations where data is often sparse and noisy. Through extensive numerical experiments, ranging from synthetic examples to complex real-world datasets such as NOAA sea-surface temperature and ERA5 wind speed, we have demonstrated the robustness and effectiveness of our approach in learning the reconstruction map under challenging conditions.
This work opens new avenues for future research, where the fusion of measure-theoretic approaches with dynamical systems theory can further enhance our ability to model, predict, and control complex systems in the presence of uncertainty [56, 8, 9, 25, 6]. Our findings not only extend the applicability of Takens’ theorem but also establish a powerful measure-theoretic framework for tackling real-world challenges in data-driven modeling and beyond.
Acknowledgements
@@
J. Botvinick-Greenhouse was supported by a fellowship award under contract FA9550-21-F-0003 through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, sponsored by the Air Force Research Laboratory (AFRL), the Office of Naval Research (ONR) and the Army Research Office (ARO). M. Oprea and Y. Yang were partially supported by the National Science Foundation through grants DMS-2409855 and by the Office of Naval Research through grant N00014-24-1-2088. R. Maulik was partially supported by U.S. Department of Energy grants, DE-FOA-0002493 and DE-FOA-0002905.
Supplemental Materials
In this supplemental text, we provide additional details to support our theoretical results in Section 3 of the main text, as well as our experimental results in Section 4 of the main text. In Section A, we provide proofs of Proposition 1, Proposition 2, Lemma 1, and Theorem 3. These results are used in the main text to derive Theorem 4 and Corollary 1, which are our paper’s central theoretical results. In Section B, we provide additional information and visualizations for our numerical experiments appearing in Section 4 of the main text.
Appendix A Proofs from Main Text

### v2

@@ 4.2.1 State Reconstruction
[…]
For the Lorenz-63 experiment, we also estimate the Lipschitz constant of the learned inverse embedding map (parameterized by a neural network) in both cases when the data is clean and noisy. The Lipschitz constant is approximated as $\sup_{z_{i}}\|\nabla\mathcal{R}_{\theta}(z_{i})\|_{2}$, where $\nabla\mathcal{R}_{\theta}$ is computed via automatic differentiation and $\{z_{i}\}_{i=1}^{10^{5}}$ are uniform samples from the unit cube, which contains the normalized training data for neural network training. We find that the Lipschitz constants of the pointwise approach when trained on clean and noisy data are 7.12 and 19.97, respectively. Moreover, the Lipschitz constants of the measure-theoretic approach are 7.56 and 7.09, respectively. The reported Lipschitz constants are the median values obtained from 10 neural network trainings each with different random initializations. While the Lipschitz constants of the measure-theoretic and pointwise approaches are comparable when the data is clean, the pointwise objective has a significantly larger Lipschitz constant when trained on noisy data. This observation is consistent with the oscillatory behavior of the pointwise reconstructions in Fig. 2(a).
4.2.2 Time-Series Prediction
While in Section 4.2.1 we studied the problem of learning the reconstruction map from the delay state to the full state, we now focus on the problem of time-series prediction. In this setting, we first form the the delay-coordinates from a partially observed scalar-valued time-series. We then aim to learn the dynamics in delay coordinates using either the pointwise (17) or measure-theoretic (18) objectives. After learning the map, we simulate the learned delay-coordinate dynamics forward in time to assess the model’s accuracy. This can be turned into a scalar time-series prediction simply by projecting the modeled delay-coordinate trajectory onto a single coordinate axis.
Figure 3: Performing time-series prediction for the partially observed Lorenz-63 system from sparse and noisy data. Fig. 3(a) shows the distribution of short-time prediction errors in delay coordinates and Fig. 3(b) shows long trajectories of the learned delay-coordinate dynamics.
In Fig. 3, we visualize results for performing time-series prediction of the scalar-valued Lorenz-63 time-series $h_{i}=x_{i}\cdot e_{1}$, where $e_{1}\in\mathbb{R}^{3}$ is the first standard unit basis vector. After forming a long delay-coordinate trajectory in dimension $m=4$ with $\tau=0.2$, we assemble a sparse and noisy dataset by randomly sampling $N=2\times 10^{3}$ initial conditions $\{\Phi(x_{i})\}_{i=1}^{N}$ in the delay-coordinates. For each chosen initial condition, we observe its final position $\{\mathcal{T}(\Phi(x_{i}))\}_{i=1}^{N}$ in delay-coordinates after a time of $\Delta t=0.1$ has elapsed. Thus, in this case, $\mathcal{T}$ corresponds to the time-$\Delta t$ (with $\Delta t=0.1$) flow map of the delay-coordinate dynamics; see (13). Based on these input-ouput training pairs, we use a constrained k-means routine to partition the initial and final positions into $K=10$ distinct empirical measures, i.e., we construct the measure-valued data $\{(\mathcal{T}\#(\Phi\#\mu_{i}),\Phi\#\mu_{i})\}_{i=1}^{K}$. We then parameterize the delay-coordinate dynamics $\mathcal{T}_{\theta}$ as the time-$\Delta t$ flow map of a neural network-parameterized velocity with node sizes $4\rightarrow 100\rightarrow 100\rightarrow 100\rightarrow 4$ and a hyperbolic tangent activation function. Throughout, all models are trained for $5\times 10^{4}$ steps using the Adam optimizer with an initial learning rate of $10^{-3}$ which decays on an exponential schedule to $10^{-5}$. We aim to learn $\mathcal{T}_{\theta}$ using both the pointwise and measure-theoretic objectives, as well as from clean data and noisy data. In the noisy case, the one-dimensional time series is corrupted by Gaussian noise with covariance $0.05$ before the delay coordinates are formed. For the measure-based learning (18) we choose $\mathfrak{D}$ as the MMD based on the kernel $k_{1}$.
The 1-step errors of the trained models, given by $|\mathcal{T}_{\theta}(\Phi(x_{i}))-\mathcal{T}(\Phi(x_{i}))|^{2}$, are visualized in Fig. 3(a) for the $\mathcal{T}_{\theta}$ trained under both the pointwise loss (17) and the measure-based loss (18). Fig. 3(b) then displays simulated delay-coordinate trajectories of both the pointwise and measure-based models after training on the sparse and noisy dataset. While each objective only trains on a 1-step loss seeking to fit $\mathcal{T}_{\theta}$, we can integrate the learned vector field over arbitrary time horizons to generate longer trajectories. As shown in Fig. 3(b), the simulated trajectory from the measure-theoretic objective appears smoother and has better visual agreement with the ground truth trajectory.
Clean data Noisy data
Pointwise $(7.70\pm 2.10)\times 10^{-4}$ $(1.42\pm 0.09)\times 10^{-1}$
Measure $(2.66\pm 0.26)\times 10^{-3}$ $(4.83\pm 0.27)\times 10^{-2}$
Table 3: Comparing the prediction error between the pointwise (17) and measure (18) approaches for forecasting the partially observed Lorenz-63 time-series. The error statistics are accumulated over 5 trials with different randomized neural network initializations.
4.3 NOAA Sea Surface Temperature Reconstruction
We now consider the problem of reconstructing the NOAA Sea Surface Temperature (SST) from partial measurement data [27]. The SST dataset consists of weekly temperature measurements sampled at a geospatial resolution of $1\degree$; see Appendix B for a visualization. Our partial observation of the full SST dataset $\{\mathbf{z}(t_{i})\}\subseteq\mathbb{R}^{44219}$ consists of the temperature time-series $\{x(t_{i})\}\subseteq\mathbb{R}$ recorded at the location $(156\degree,40\degree)$. Our goal is to use the delay state corresponding to $\{x(t_{i})\}$ to learn a reconstruction map for the full state $\{\mathbf{z}(t_{i})\}$. That is, we aim to learn $\Phi^{-1}$ which maps $\{x(t_{i})\}$ into $\{\mathbf{z}(t_{i}))\}$. The success of the learned reconstruction map is then evaluated by applying it to unseen testing data in delay coordinates. A similar problem is considered in [51, 50].
@@ 4.3 NOAA Sea Surface Temperature Reconstruction
The dataset is partitioned into training and testing sets, where the training set contains 355 snapshots ($\approx 5$ years) and the testing set contains 1415 snapshots $(\approx 27$ years). Based on the time-series $\{x(t_{i})\}$, we select a time delay of $\tau=12$ (weeks) and an embedding dimension $m=7$. To reduce the computational cost in learning, similar to [50, 51], we subtract off the temporal mean of the SST dataset and parameterize the full state of the system by the first $N_{\texttt{POD}}$ time-varying POD coefficients, $\{\alpha_{k}(t)\}_{k=1}^{N_{\texttt{POD}}}$, which are obtained via the method of snapshots; see [50, Section 3.1]. That is, we perform the model reduction
@@
For this problem, all training samples are normalized via an affine transformation, such that the $L^{\infty}$ norm of the data vectors is at most 1. We train the neural network parameterization $\mathcal{R}_{\theta}$ using both the pointwise (see Eqn. (14)) and measure-based (see Eqn. (15)) approaches. The pointwise approach seeks to directly enforce the relationship (20) through the mean-squared error, whereas the measure-based approach partitions the delay state into clusters and aims to push forward the empirical measure in each cluster into the corresponding measure in the POD space. We use a constrained k-means routine to evenly partition the delay state into $5$ clusters. A visualization of the 5 clusters can be found in Appendix B. We evaluate the performance of the learned $\mathcal{R}_{\theta}$ by reconstructing the time-varying POD coefficients $\{\alpha_{k}(t_{i})\}$ for the testing set, which are then used to generate the full state $\{\mathbf{z}(t_{i})\}$ according to (19). More specifically, we reconstruct the $\{\alpha_{k}(t_{i})\}$ by applying the trained model $\mathcal{R}_{\theta}$ to the delay state based on the time-series $\{x(t_{i})\}.$
We note that the challenge of learning the reconstruction map $\mathcal{R}_{\theta}:\mathbb{R}^{7}\to\mathbb{R}^{200}$ is exacerbated by the sparsity of the available training data. Specifically, we aim to learn a $200$-dimensional map using only $355$ training examples. Due to this data sparsity, we utilize MMD based on the Gaussian kernel $k_{2}(x,y)=\exp(-\|x-y\|_{2}^{2}/2\sigma^{2})$ with $\sigma=3$. The choice of a relatively large $\sigma$ acts as a form of regularization, helping to mitigate overfitting when dealing with sparse samples [65].
@@
Table 4: Reconstruction mean squared error (MSE) for the SST dataset after various numbers of training iterations. The measure-based approach (see Eqn. (15)) is less prone to overfitting than the pointwise approach (see Eqn. (14)).
In Fig. 4, we visualize the pointwise and measure-based reconstructions of the SST example at testing weeks 425, 550, 675, and 800 after training both models for 25,00025{,}000 steps. One can observe that the measure-based results align more closely with the ground-truth snapshots, while the pointwise reconstructions exhibit numerous nonphysical oscillations.
Figure 4: Visual comparison of the pointwise (see (14)) and measure-based (see (15)) approaches to reconstructing the SST dataset at the testing weeks 425, 550, 675, and 800. The left column features the ground truth snapshot, the middle column shows the measure-based reconstruction, and the right column shows the pointwise reconstruction.
@@
Table 5: Testing reconstruction mean squared error (MSE) for the ERA5 wind speed dataset after various iterations during training. For each loss function, the experiment was repeated 55 times to approximate the mean and standard deviation for the reconstruction error.
@@ 4.4 ERA5 Wind Field Reconstruction
Our final test reconstructs a portion of the ERA5 wind speed dataset using partial measurement data [28]. Here, we will again aim to learn the full-state reconstruction map $\Phi^{-1}$ using training data from both the full state and a delay state corresponding to a partial observable. We consider the ERA5 wind dataset sampled at a geopotential height of Z500, restricted to latitudes between $180\degree$ and $105\degree$ and longitudes between $90\degree$ and $15\degree$. The dataset is coarsened in both space and time, with measurements taken on approximately a $4\degree$ spatial grid with dimensions $20\times 20$, and separated by $6$ hours in time. For training, we use $2\times 10^{3}$ randomly sampled snapshots of both the $u$ and $v$ wind speed components (longitudinal and latitudinal, respectively) between 2000 and 2007. For testing, we consider the subsequent $5\times 10^{3}$ consecutive snapshots.
@@
In (21), the time-delay and embedding dimension are heuristically chosen as $\tau=6$ (hours) and $m=4$. In (22), the indices $\{n_{k}\}_{k=1}^{30}$ correspond to 30 randomly sampled spatial locations at which we observe the longitudinal and latitudinal components of the wind speed. While it was possible to obtain a reasonable model for the full state of the NOAA SST dataset from a single observable, the wind speed dynamics considered here exhibit significantly more complex spatio-temporal behavior. Thus, we find it necessary to use measurements from several spatial locations to perform accurate state reconstruction. The number 30 was also chosen heuristically. The full longitudinal wind speed vector is $\mathbf{u}(t_{i})=\left(u_{1}(t_{i}),\dots,u_{400}(t_{i})\right)\in\mathbb{R}^{400}$ representing values at the $20\times 20$ spatial grid, while the full latitudinal wind speed vector is $\mathbf{v}(t_{i})=\left(v_{1}(t_{i}),\dots,v_{400}(t_{i})\right)\in\mathbb{R}^{400}.$ The combined full state can then be written $\mathbf{z}(t_{i})=(\mathbf{u}(t_{i}),\mathbf{v}(t_{i}))\in\mathbb{R}^{800}$, which is precisely the quantity we seek to predict from the delay state $\mathbf{y}(t_{i})$; see (21).
We train models to approximate the state reconstruction map $\mathbf{y}(t_{i})\mapsto\mathbf{z}(t_{i})$ using both the pointwise reconstruction loss (14) and the measure-based reconstruction loss (15). Throughout, we normalize the training data by subtracting the temporal mean and ensuring that each feature has unit variance. Recall that we randomly sample $2\times 10^{3}$ times $t_{i}$ at which we observe both the delay state $\mathbf{y}(t_{i})$ and the full state $\mathbf{z}(t_{i})$, allowing us to form the paired data $\{(\mathbf{z}(t_{i}),\mathbf{y}(t_{i}))\}_{i=1}^{2000}.$ Thus, when learning the reconstruction map according to the pointwise loss (14), we directly enforce the relationship $\mathbf{y}(t_{i})\mapsto\mathbf{z}(t_{i})$, for $1\leq i\leq 2\times 10^{3}$, via the mean squared error.
In contrast, when learning the reconstruction map using the measure-based loss (15), we transform the paired data $\{(\mathbf{z}(t_{i}),\mathbf{y}(t_{i}))\}_{i=1}^{2000}$ into probability measures and aim to find a suitable pushforward map between the probability measures in the time-delayed state space and those in the full state space. We first apply a constrained k-means clustering algorithm to partition the time-delayed samples $\{\mathbf{y}(t_{i})\}_{i=1}^{2000}\subseteq\mathbb{R}^{240}$ into 20 distinct clusters, each containing 100 samples in $\mathbb{R}^{240}$. In doing so, we obtain 20 pairs of probability distributions. During training, we then learn a model that pushes forward each measure in the time-delay coordinate system to the corresponding measure in the full state space. We evaluate the measure-based approach (15) using both the Energy MMD loss (with the kernel $k_{1}$) and the Gaussian MMD loss (with a Gaussian kernel $k_{2}$ and $\sigma=25$).
@@
It is worth noting that choosing a relatively large bandwidth for the Gaussian MMD acts as additional regularization, helping to prevent overfitting to the locations of the training samples. For longer training times, the Gaussian MMD outperforms the Energy MMD in measure-based reconstruction. The reconstructions based on the Gaussian MMD are smoother than the pointwise reconstructions and more closely match the ground truth; a visualization can be found in Appendix B. In Fig. 5, we show the average spatial distribution of the reconstruction error for the Gaussian MMD and pointwise approaches.
@@ 5 Conclusion
This work presents a measure-theoretic variant of the classical Takens embedding theorem. While the original theory focuses on the embedding property of individual time-delayed trajectories, the proposed measure-theoretic generalization shifts the focus to probability distributions over the state space. Through Corollary 9, our main theoretical contribution, we have demonstrated that the embedding also holds within the space of probability distributions under the pushforward action of the time-delay map. This generalization is obtained by combining the classical Takens’ theorem with geometric tools from the theory of optimal transport.
@@
Building on these theoretical foundations, we devised a measure-theoretic computational routine for learning the inverse embedding map, also known as the full-state reconstruction map. This method enables the reconstruction of an entire high-dimensional dynamical system state from the time lags of a single observable. Our approach involves partitioning the observed trajectory in time-delay coordinates using a k-means clustering routine, where each cluster represents discrete probability measures. During training, we ensure that the corresponding discrete measures in the reconstruction space match the pushforward distributions of the measures in the time-delay coordinates. This training scheme represents a relaxation of classical pointwise matching, allowing for greater tolerance of pointwise errors in the final reconstruction based on the scale of each discrete measure in the reconstruction space.
While pointwise matching can be accurate under idealized, noise-free and densely sampled conditions, such settings are uncommon in practice. In scenarios with sparse or noisy observations, our distributional training can offer advantages. On synthetic tests and two geophysical datasets (NOAA sea-surface temperature and ERA5 wind speed) we observe stable reconstructions and competitive accuracy for the learned reconstruction map under challenging conditions.
These results suggest several directions for further work, including integrating measure-theoretic tools with dynamical systems to model, predict, and control systems under uncertainty [66, 5, 67, 68, 69]. They extend the scope of Takens-style embedding results and contribute to a measure-theoretic perspective on data-driven modeling. We will examine the stability of measure-based loss in future work and position it alongside established results for pointwise learning. This problem is nontrivial: stability depends jointly on the function approximation class, the magnitude and structure of noise, and the choice of distance or divergence on probability measures used in the optimization framework.
[…]
@@
Acknowledgements J. Botvinick-Greenhouse was supported by a fellowship award under contract FA9550-21-F-0003 through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, sponsored by the Air Force Research Laboratory (AFRL), the Office of Naval Research (ONR) and the Army Research Office (ARO). M. Oprea and Y. Yang were partially supported by the National Science Foundation through grants DMS-2409855 and by the Office of Naval Research through grant N00014-24-1-2088. R. Maulik was partially supported by U.S. Department of Energy grants, DE-FOA-0002493 and DE-FOA-0002905.
Declarations
  • Funding: J. Botvinick-Greenhouse was supported by a fellowship award under contract FA9550-21-F-0003 through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, sponsored by the Air Force Research Laboratory (AFRL), the Office of Naval Research (ONR) and the Army Research Office (ARO). M. Oprea and Y. Yang were partially supported by the National Science Foundation through grants DMS-2409855 and by the Office of Naval Research through grant N00014-24-1-2088. R. Maulik was partially supported by U.S. Department of Energy grants, DE-FOA-0002493 and DE-FOA-0002905.
  • Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
  • Ethics approval and consent to participate: This does not apply to this article.
  • Consent for publication: All authors consent to publish this article.
  • Data availability: The data used in this article is available to download from the open-source repositories.
  • Materials availability: All materials related to this work are available to download from open-source domains.
  • Code availability: the code is available at https://github.com/jrbotvinick/Measure-Theoretic-Time-Delay-Embedding.
  • Author contribution: J. B.-G., M. O., R. M. and Y. Y. designed the research, J. B.-G. and M. O. performed research; J. B.-G. , M. O. and Y. Y. contributed new reagents or analytic tools; J. B.-G. and R. M. analyzed data; J. B.-G., M. O. and Y. Y. wrote the paper; R. M. reviewed the paper.
Appendix A Proofs from the Main Text